# Polarons and Exciton-Polarons in Two-Dimensional Polar Materials

### v1

@@ Abstract
Exciton-polaron. When an exciton forms in a 2D polar crystal, both electron and hole couple to LO phonons. The resulting composite quasiparticle can be viewed as an exciton dressed by a phonon cloud, i.e., an exciton-polaron. The Hamiltonian of an electron-hole pair interacting with LO phonons is
@@
$+\sum_{\mathbf{k}}\left(V_{k}e^{i\mathbf{k}\cdot\mathbf{r}_{e}}\hat{a}_{\mathbf{k}}+V_{k}^{*}e^{-i\mathbf{k}\cdot\mathbf{r}_{e}}\hat{a}^{\dagger}_{\mathbf{k}}\right)$
$-\sum_{\mathbf{k}}\left(V_{k}e^{i\mathbf{k}\cdot\mathbf{r}_{h}}\hat{a}_{\mathbf{k}}+V_{k}^{*}e^{-i\mathbf{k}\cdot\mathbf{r}_{h}}\hat{a}^{\dagger}_{\mathbf{k}}\right),$ (11)
where $V_{\rm KR}$ is the bare 2D Coulomb (Keldysh-Rytova) potential, reading as [48, 49, 50]
@@
with $H_{0}$, $Y_{0}$ denoting the Struve function, and the Bessel function of the second kind. We transform to center of center-of-mass $\mathbf{R}=(m_{e}\mathbf{r}_{e}+m_{h}\mathbf{r}_{h})/M$, and relative coordinates $\mathbf{r}=\mathbf{r}_{e}-\mathbf{r}_{h}$, where $M=m_{e}+m_{h}$ is the exciton total mass. We then adopt a variational approach developed for bulk exciton-polarons [18]. Particularly, we apply a unitary transformation $\hat{S}_{\rm X}=\exp\left(i\left(\mathbf{Q}-\sum_{k}\mathbf{k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\right)\cdot\mathbf{R}\right)$ which eliminates the center-of-mass coordinate. Here $\hbar\mathbf{Q}$ denotes the total momentum of exciton. Similar to the case of single polaron, we introduce the mean value of displacement field $F_{\mathbf{k}}(\mathbf{r})$, and apply a unitary transformation with operator $\hat{U}_{\rm X}=\exp\left(\sum_{k}\left(\hat{a}_{\mathbf{k}}F^{*}_{\mathbf{k}}(\mathbf{r})-\hat{a}_{\mathbf{k}}^{\dagger}F_{\mathbf{k}}(\mathbf{r})\right)\right)$. Assuming again that the resulting wave function corresponds to a phonon ground-state $\hat{a}_{\mathbf{k}}|\psi_{\rm X}\rangle=0$, and setting the center-of-mas momentum $\hbar\mathbf{Q}=0$, we get the following energy functional (see SM):
$E_{\rm X}$ $=\left\langle\psi_{\rm X}\left|\hat{\mathbf{p}}^{2}/(2\mu)+V_{\rm KR}(r,r_{\infty})+W_{\rm X}(\mathbf{r})\right|\psi_{\rm X}\right\rangle,$ (13)
where $\mu=m_{e}m_{h}/(m_{e}+m_{h})$ is the exciton reduced mass,
@@
$W_{\rm X}(\mathbf{r})=\frac{\hbar^{2}}{2\mu}\sum_{k}\nabla F^{*}_{\mathbf{k}}\cdot\nabla F_{\mathbf{k}}+\sum_{k}\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2M}\right)|F_{\mathbf{k}}|^{2}\\ -\sum_{k}\left(V_{k}\rho_{\mathbf{k}}F_{\mathbf{k}}+V_{k}^{*}\rho_{\mathbf{k}}^{*}F_{\mathbf{k}}^{*}\right),$ (14)
@@
and $\rho_{\mathbf{k}}=\exp\left(im_{h}\mathbf{k}\cdot\mathbf{r}/M\right)-\exp\left(-im_{e}\mathbf{k}\cdot\mathbf{r}/M\right)$. Unlike the case of single polarons, here the presence of the relative coordinate and relative momentum operator $\hat{\mathbf{p}}$ couples the displacement field $F_{\mathbf{k}}(\mathbf{r})$ and the exciton wave function $\psi_{\rm X}$. For the sake of simplicity, we ignore this mixing, and apply the variational minimization $\partial W_{\rm X}/\partial F_{\mathbf{k}}=0$. The result reads as follows (see SM):
$E_{\rm X}^{\rm b}$ $=\left\langle\psi_{\rm X}\left|\hat{\mathbf{p}}^{2}/(2\mu^{*})+V_{\rm eff}(r)\right|\psi_{\rm X}\right\rangle.$ (15)
Here we defined the exciton binding energy as $E_{\rm X}^{\rm b}=E_{\rm X}+E^{\rm b}_{e}+E^{\rm b}_{h}$, and phenomenologically included the mass renormalizations: $\mu^{*}=m_{e}^{*}m_{h}^{*}/(m_{e}^{*}+m_{h}^{*})$. Such renormalization naturally emerges in the full treatment [19]. The effective Coulomb potential can be written as
$V_{\rm eff}^{\rm ap}(r)$ $\approx V_{\rm KR}(r,r_{0})$
@@
$-E_{0}\left[\frac{m_{h}}{\Delta m}h_{\tilde{\sigma}_{h,t}}\left(\frac{\varepsilon r}{r_{\infty}}\right)-\frac{m_{e}}{\Delta m}h_{\tilde{\sigma}_{e,t}}\left(\frac{\varepsilon r}{r_{\infty}}\right)\right],$ (16)
where $\Delta m=m_{h}-m_{e}$, and
$h_{\nu}(x)=\frac{\sigma_{0}-1}{2(1+\nu^{2})(\sigma_{0}^{2}+\nu^{2})}\left(2(\nu^{2}-\sigma_{0})K_{0}\left(x/\nu\right)\right.$
$\left.+(1+\sigma_{0})\alpha\pi\left[I_{0}\left(x/\nu\right)-L_{0}\left(x/\nu\right)\right]\right)$ (17)
@@
is the function expressing the phonon-induced corrections. Here $I_{0}$, $K_{0}$ are the modified Bessel functions of the first and the second kind, respectively, and $L_{0}$ is the modified Struve function. The Eq. (16) was derived in an approximation $\tilde{\sigma}_{e,{\rm t}}^{2},\tilde{\sigma}_{h,{\rm t}}^{2}\ll 1$, and the total expression is given in SM.
The modified Coulomb potential given by Eq. (16) is the 2D counterpart of Pollmann-Büttner potential, characterizing the bulk exciton-polarons [18]. Fig. 3 (a), (b) illustrate the coordinate dependence of effective potential for hBN and HfSe2 monolayers, respectively. Near the origin $V_{\rm eff}\rightarrow V_{\rm KR}(r,r_{\infty})$, whereas for the long-range tail $V_{\rm eff}\rightarrow V_{\rm KR}(r,r_{0})$, analogous to the case of bulk polarons [20]. The polaronic modification of the effective potential is more pronounced for HfSe2 monolayer shown in Fig. 3 (b), which is due to strong polar effects characterized by the condition $\sigma_{0}\gg 1$.
Using the obtained effective potential, we evaluate the exciton-polaron binding energy, defined by Eq. (15). Figures 4 (a), (b) illustrate the exciton wave functions in hBN and HfSe2 monolayers, respectively. The dotted curves correspond to bare exciton states, discarding the polaron effects. The drastic reshaping of the wave function in the HfSe2 monolayer is again the manifestation of strong polaronic effects. This also translates to the values of exciton binding energy (314.9 meV vs 127.5 meV for bare exciton). Here the binding energy of the bare exciton $E_{\rm X}^{0}$ is calculated with the statically screened potential $V_{\rm KR}(r,r_{0})$. For hBN monolayer the polaron effects are less pronounced, and the exciton-polaron binding energy is 1869.2 meV (1616 meV for bare exciton). Fig. 4 (c) summarizes both the single polaron and exciton-polaron binding energy data for a range of polar atomic monolayers. Here $\Delta E_{\rm X}^{\rm b}=E_{\rm X}^{\rm b}-E_{\rm X}^{0}$ characterizes the modulation of exciton binding energy stemming from polaron effects.
Conclusion. We developed a variational theory of polaron and exciton-polaron formation in 2D polar materials. Our macroscopic model allows to derive an analytical expression for LO phonon dispersion via simple electrostatic considerations, which reproduces the results of first-principal calculations [52]. Variational treatment of electron-phonon coupling problem results in semi-analytical expressions for the binding energy and mass renormalization of Frohlich polaron, which essentially differs from its bulk counterpart [6]. The extension of variational method for the exciton-polaron problem reveals the modulation of pure 2D electron-hole Coulomb interaction due to the coupling of charge carriers to phonons, in line with the case of bulk exciton-polarons [18]. The framework of pure 2D polarons presented here can be readily expanded to account for the finite thickness, characteristic to layered hybrid perovskites [66].
@@ I.1 Lagrangian density
The Lagrangian density of the system is given by:
@@ II LO phonon Hamiltonian
We recall that $\bm{\xi}(\mathbf{r})=\sqrt{MN/\Sigma}\mathbf{S}(\mathbf{r})$, and introduce a plane-wave expansion
@@
$\frac{\omega_{\rm t}^{2}}{2}\int\bm{\xi}^{2}(\mathbf{r})\delta(z){\rm d^{2}}r{\rm d}z$ $=\frac{\omega_{\rm t}^{2}}{2}\frac{MN}{\Sigma N}\int\sum_{k}\sum_{k^{\prime}}e^{i(\mathbf{k}+\mathbf{k}^{\prime})\cdot\mathbf{r}}\mathbf{X}_{\mathbf{k}}\cdot\mathbf{X}_{\mathbf{k}^{\prime}}{\rm d^{2}}r=\frac{M\omega_{\rm t}^{2}}{2}\sum_{k}\mathbf{X}_{\mathbf{k}}\cdot\mathbf{X}_{-\mathbf{k}},$ (S33)
$\frac{1}{2}\int\dot{\bm{\xi}}^{2}(\mathbf{r})\delta(z){\rm d^{2}}r{\rm d}z$ $=\frac{1}{2}\frac{MN}{\Sigma N}\int\sum_{k}\sum_{k^{\prime}}e^{i(\mathbf{k}+\mathbf{k}^{\prime})\cdot\mathbf{r}}\dot{\mathbf{X}}_{\mathbf{k}}\cdot\dot{\mathbf{X}}_{\mathbf{k}^{\prime}}{\rm d^{2}}r=\frac{M}{2}\sum_{k}\dot{\mathbf{X}}_{\mathbf{k}}\cdot\dot{\mathbf{X}}_{-\mathbf{k}}=\frac{1}{2M}\sum_{k}\mathbf{P}_{\mathbf{k}}\cdot\mathbf{P}_{-\mathbf{k}},$ (S34)
where $\mathbf{P}_{\mathbf{k}}=M\dot{\mathbf{X}}_{\mathbf{k}}$. Noting that $\bm{\xi}_{\mathbf{k}}=\mathbf{X}_{\mathbf{k}}\sqrt{\Sigma M}/(2\pi)^{2}$, we get
@@
$=\frac{\gamma^{2}}{2\varepsilon_{0}}\sqrt{\frac{MN}{\Sigma}}\frac{1}{\sqrt{N}}\int\sum_{k}e^{i\mathbf{k}\cdot\mathbf{r}}\frac{q}{\varepsilon+r_{\infty}q}e^{i\mathbf{q}\cdot\mathbf{r}}\frac{\sqrt{\Sigma M}}{(2\pi)^{2}}\mathbf{X}_{\mathbf{q}}^{||}\cdot\mathbf{X}_{\mathbf{k}}{\rm d^{2}}r{\rm d^{2}}q$
@@
$=\frac{\alpha_{\infty}}{2}\frac{\gamma^{2}}{(2\varepsilon_{0})^{2}}\int\frac{q}{\varepsilon+r_{\infty}q}\frac{k}{\varepsilon+r_{\infty}k}\delta(\mathbf{q}+\mathbf{k})(2\pi)^{2}\frac{\Sigma M}{(2\pi)^{4}}\mathbf{X}_{\mathbf{q}}^{||}\cdot\mathbf{X}_{\mathbf{k}}^{||}{\rm d^{2}}q{\rm d^{2}}k$
@@ III Electron-phonon coupling
$\bm{\xi}^{||}(\mathbf{r})$ $=\sqrt{\frac{MN}{\Sigma}}\mathbf{S}^{||}(\mathbf{r})=\sum_{k}\sqrt{\frac{\hbar}{2\Sigma\omega_{{\rm l},k}}}\left(e^{i\mathbf{k}\cdot\mathbf{r}}a_{\mathbf{k}}+e^{-i\mathbf{k}\cdot\mathbf{r}}a_{\mathbf{k}}^{\dagger}\right)\mathbf{e}_{\mathbf{k}}.$ (S43)
@@
$\bm{\xi}_{\mathbf{q}}^{||}=\frac{\sqrt{\Sigma M}}{(2\pi)^{2}}\mathbf{X}_{\mathbf{q}}^{||}=\sqrt{\frac{\hbar\Sigma}{2\omega_{{\rm l},q}}}\frac{1}{(2\pi)^{2}}\left(a_{\mathbf{q}}+a_{-\mathbf{q}}^{\dagger}\right)\mathbf{e}_{\mathbf{q}},$ (S45)
@@
$\varphi(\mathbf{r},0)$ $=-\frac{i\gamma}{2\varepsilon_{0}}\int e^{i\mathbf{q}\cdot\mathbf{r}}\frac{1}{q}\frac{1}{\varepsilon+r_{\infty}q}\mathbf{q}\cdot\sqrt{\frac{\hbar\Sigma}{2\omega_{{\rm l},q}}}\frac{1}{(2\pi)^{2}}\left(a_{\mathbf{q}}+a_{-\mathbf{q}}^{\dagger}\right)\mathbf{e}_{\mathbf{q}}{\rm d}^{2}q$
$=-i\omega_{\rm t}\sum_{q}\sqrt{\frac{\hbar(r_{0}-r_{\infty})}{4\varepsilon_{0}\Sigma\omega_{{\rm l},q}}}\frac{1}{\varepsilon+r_{\infty}q}\left(a_{\mathbf{q}}e^{i\mathbf{q}\cdot\mathbf{r}}-a_{\mathbf{q}}^{\dagger}e^{-i\mathbf{q}\cdot\mathbf{r}}\right).$ (S46)

### v2

@@ Abstract
Exciton-polaron. When an exciton forms in a 2D polar crystal, both electron and hole couple to LO phonons. The resulting composite quasiparticle, an exciton dressed by a phonon cloud, is referred to as an exciton-polaron. The Hamiltonian of an electron-hole pair interacting with LO phonons is
@@
$+\sum_{\mathbf{k}}\left(V_{k}\hat{a}_{\mathbf{k}}\left(e^{i\mathbf{k}\cdot\mathbf{r}_{e}}-e^{i\mathbf{k}\cdot\mathbf{r}_{h}}\right)+{\rm h.c.}\right),$ (11)
where $V_{\rm KR}$ is the "bare" 2D Coulomb (Keldysh-Rytova) potential [50, 51, 52] with a purely electronic screening,
@@
where $H_{0}$, $Y_{0}$ are the Struve function, and the Bessel function of the second kind. We then follow a variational approach developed for 3D exciton-polarons [18, 19], developed as a generalization of the single polaron LLP theory [6]. The main feature of excitonic generalization is that the variational variable $F_{\mathbf{k}}(\mathbf{r})$, introduced through a shift transformation $\hat{U}^{-1}\hat{a}_{\mathbf{k}}\hat{U}=\hat{a}_{\mathbf{k}}+F_{\mathbf{k}}(\mathbf{r})$, becomes a function of the relative coordinate $\mathbf{r}=\mathbf{r}_{e}-\mathbf{r}_{h}$. The variational energy, which also depends on the exciton wave function $\psi_{X}(\mathbf{r})$, takes the form (see [19] and SM):
$E_{\rm X}$ $=\left\langle\psi_{\rm X}\left|\hat{\mathbf{p}}^{2}/2\mu+V_{\rm KR}(r,r_{\infty})+W_{\rm X}(\mathbf{r})\right|\psi_{\rm X}\right\rangle,$ (13)
where $\mu=m_{e}m_{h}/M$, $M=m_{e}+m_{h}$ are the exciton reduced and total mass, and
@@
$W_{\rm X}(\mathbf{r})=\sum_{k}\Big{\{}\frac{\hbar^{2}}{2\mu}|\nabla F_{\mathbf{k}}|^{2}+\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2M}\right)|F_{\mathbf{k}}|^{2}\\ -\left[V_{k}\left(e^{i\frac{m_{e}}{M}\mathbf{k}\cdot\mathbf{r}}-e^{i\frac{m_{h}}{M}\mathbf{k}\cdot\mathbf{r}}\right)F_{\mathbf{k}}+{\rm c.c.}\right]\Big{\}},$ (14)
@@
Strictly speaking, the presence of the relative momentum operator $\hat{\mathbf{p}}$ couples the displacement field $F_{\mathbf{k}}(\mathbf{r})$ and the exciton wave function $\psi_{\rm X}$. For simplicity, we ignore this mixing, and apply the variational minimization as $\delta W_{\rm X}/\delta F_{\mathbf{k}}=0$, which determines an effective electron-hole interaction potential $V_{\rm eff}(\mathbf{r})$ dressed by LO phonons. As a result, the exciton energy is computed as follows,
$E_{\rm X}$ $=-E^{\rm b}_{e}-E^{\rm b}_{h}+\left\langle\psi_{\rm X}\left|\hat{\mathbf{p}}^{2}/2\mu^{*}+V_{\rm eff}(r)\right|\psi_{\rm X}\right\rangle,$ (15)
where the negative of the last term defines the binding energy $E_{\rm X}^{\rm b}$ of polaron-exciton, and $\mu^{*}=m_{e}^{*}m_{h}^{*}/(m_{e}^{*}+m_{h}^{*})$ is the renormalized mass  [^5] Such renormalization of the reduced mass naturally emerges in a more accurate treatment of the variational problem [19]. The exact form of $V_{\rm eff}(r)$ can be found in SM, Sec. V. In the practically relevant case of $\tilde{\sigma}_{e,{\rm t}}^{2},\tilde{\sigma}_{h,{\rm t}}^{2}<1$, it simplifies as follows,
@@
$V_{\rm eff}^{\rm ap}(r)\approx V_{\rm KR}(r,r_{0})\\ +E_{0}\left[\frac{m_{h}}{\Delta m}\Phi_{\tilde{\sigma}_{h,t}}\!\!\left(\frac{\varepsilon r}{r_{\infty}}\right)-\frac{m_{e}}{\Delta m}\Phi_{\tilde{\sigma}_{e,t}}\!\!\left(\frac{\varepsilon r}{r_{\infty}}\right)\right],$ (16)
where $\Delta m=m_{h}-m_{e}$, and the function
$\Phi_{\nu}(x)=\frac{\sigma_{0}-1}{2(1+\nu^{2})(\sigma_{0}^{2}+\nu^{2})}\big{\{}2(\nu^{2}-\sigma_{0})K_{0}\left(x/\nu\right)$
$+\pi\nu(1+\sigma_{0})\left[I_{0}\left(x/\nu\right)-L_{0}\left(x/\nu\right)\right]\big{\}}$ (17)
@@
encodes phonon-induced corrections. Here $I_{0}$, $K_{0}$ are the modified Bessel functions of the first and the second kind, respectively, and $L_{0}$ is the modified Struve function. Equation (16) represents a 2D counterpart of the Pollmann-Büttner potential, characterizing exciton-polarons in 3D [18, 20].
The potential $V_{\rm eff}(r)$ for hBN and HfSe2 monolayers, is shown in Fig. 3 (a) and (b), respectively. Apparently, the polaron effects are more pronounced in HfSe2 due to stronger coupling with $\sigma_{0}\gg 1$. A generic feature of $V_{\rm eff}(r)$ is that it interpolates between the "bare", electron-only screened KR potential, $V_{\rm eff}(r)\rightarrow V_{\rm KR}(r,r_{\infty})$, near the origin, and the fully screened potential, $V_{\rm eff}(r)\rightarrow V_{\rm KR}(r,r_{0})$, at large distances. This translates into exciton properties, defined by Eq. (15). In particular, in the strongly coupled case of HfSe2 the exciton-polaron binding energy is strongly reduced to 306.9 meV from $E_{\rm X}^{\infty}=421.4$ meV for the bare exciton; see Table 4. However, this value is still significantly higher than $E_{\rm X}^{0}=127.5$ meV obtained naively using the statically screened KR potential $V_{\rm KR}(r,r_{0})$. In hBN monolayer the polaron screening effects are less pronounced, but the reduction of the binding energy is still sizable, 1865 meV vs 2007 meV for the bare exciton. The data for polaron and exciton-polaron binding in a range of materials are summarized in Fig. 4.
Conclusion. In conclusion, we developed a macroscopic theory of LO phonons, polarons and exciton-polarons in 2D polar materials. Our simple analytic consideration, involving only intrinsic 2D polarizabilities, perfectly agrees with the results of first-principal calculations and recent experiments [53, 56]. The variational treatment of the polaron problem results in a semi-analytical expression that defines an upper bound for the energy of Fröhlich polaron, which is strongly different from its bulk counterpart [6]. The extension of the variational approach to the exciton-polaron problem reveals a modification of the electron-hole interaction by LO phonons, which leads to a dramatic change of the exciton binding energy. Our general framework opens a way for the application of more advanced mathematical techniques to study polaron effects in 2D materials. Apparently, the present purely 2D theory can be readily extended to account for the layered structure of hybrid perovskites [65].
Acknowledgments. The research is supported by the Ministry of Science and Higher Education of the Russian Federation (Goszadaniye) Project No. FSMG-2023-0011. V.S. acknowledges the support of “Basis” Foundation (Project No. 22-1-3-43-1). The work of A.K. is supported by the Icelandic Research Fund (Rannsóknasjóður, Grant No. 2410550). IVT acknowledges support from the Spanish MCIN/AEI/ 10.13039/501100011033 through the project PID2023-148225NB-C32, and the Basque Government (Grant No. IT1453-22).
@@ I.1 Lagrangian density
Let us define the mass-weighted displacement field $\bm{\xi}=\sqrt{nM}\mathbf{S}$, where $\mathbf{S}=\mathbf{U}_{+}-\mathbf{U}_{-}$ is the relative displacement of positively and negatively charged ions, $n=N/A$ is the sheet density of unit cells, and $M=m_{+}m_{-}/(m_{+}+m_{-})$ is the reduced ionic mass. The Lagrangian density of the system is given by:
@@ II LO phonon Hamiltonian
We recall that $\bm{\xi}(\mathbf{r})=\sqrt{MN/A}\mathbf{S}(\mathbf{r})$, and introduce a plane-wave expansion
@@
$\frac{\omega_{\rm t}^{2}}{2}\int\bm{\xi}^{2}(\mathbf{r})\delta(z){\rm d^{2}}r{\rm d}z$ $=\frac{\omega_{\rm t}^{2}}{2}\frac{MN}{AN}\int\sum_{k}\sum_{k^{\prime}}e^{i(\mathbf{k}+\mathbf{k}^{\prime})\cdot\mathbf{r}}\mathbf{X}_{\mathbf{k}}\cdot\mathbf{X}_{\mathbf{k}^{\prime}}{\rm d^{2}}r=\frac{M\omega_{\rm t}^{2}}{2}\sum_{k}\mathbf{X}_{\mathbf{k}}\cdot\mathbf{X}_{-\mathbf{k}},$ (S33)
$\frac{1}{2}\int\dot{\bm{\xi}}^{2}(\mathbf{r})\delta(z){\rm d^{2}}r{\rm d}z$ $=\frac{1}{2}\frac{MN}{AN}\int\sum_{k}\sum_{k^{\prime}}e^{i(\mathbf{k}+\mathbf{k}^{\prime})\cdot\mathbf{r}}\dot{\mathbf{X}}_{\mathbf{k}}\cdot\dot{\mathbf{X}}_{\mathbf{k}^{\prime}}{\rm d^{2}}r=\frac{M}{2}\sum_{k}\dot{\mathbf{X}}_{\mathbf{k}}\cdot\dot{\mathbf{X}}_{-\mathbf{k}}=\frac{1}{2M}\sum_{k}\mathbf{P}_{\mathbf{k}}\cdot\mathbf{P}_{-\mathbf{k}},$ (S34)
where $\mathbf{P}_{\mathbf{k}}=M\dot{\mathbf{X}}_{\mathbf{k}}$. Noting that $\bm{\xi}_{\mathbf{k}}=\mathbf{X}_{\mathbf{k}}\sqrt{AM}/(2\pi)^{2}$, we get
@@
$=\frac{\gamma^{2}}{2\varepsilon_{0}}\sqrt{\frac{MN}{A}}\frac{1}{\sqrt{N}}\int\sum_{k}e^{i\mathbf{k}\cdot\mathbf{r}}\frac{q}{\varepsilon+r_{\infty}q}e^{i\mathbf{q}\cdot\mathbf{r}}\frac{\sqrt{AM}}{(2\pi)^{2}}\mathbf{X}_{\mathbf{q}}^{||}\cdot\mathbf{X}_{\mathbf{k}}{\rm d^{2}}r{\rm d^{2}}q$
@@
$=\frac{\alpha_{\infty}}{2}\frac{\gamma^{2}}{(2\varepsilon_{0})^{2}}\int\frac{q}{\varepsilon+r_{\infty}q}\frac{k}{\varepsilon+r_{\infty}k}\delta(\mathbf{q}+\mathbf{k})(2\pi)^{2}\frac{AM}{(2\pi)^{4}}\mathbf{X}_{\mathbf{q}}^{||}\cdot\mathbf{X}_{\mathbf{k}}^{||}{\rm d^{2}}q{\rm d^{2}}k$
@@ III Electron-phonon coupling
$\bm{\xi}^{||}(\mathbf{r})$ $=\sqrt{\frac{MN}{A}}\mathbf{S}^{||}(\mathbf{r})=\sum_{k}\sqrt{\frac{\hbar}{2A\omega_{{\rm l},k}}}\left(e^{i\mathbf{k}\cdot\mathbf{r}}a_{\mathbf{k}}+e^{-i\mathbf{k}\cdot\mathbf{r}}a_{\mathbf{k}}^{\dagger}\right)\mathbf{e}_{\mathbf{k}}.$ (S43)
@@
$\bm{\xi}_{\mathbf{q}}^{||}=\frac{\sqrt{AM}}{(2\pi)^{2}}\mathbf{X}_{\mathbf{q}}^{||}=\sqrt{\frac{\hbar A}{2\omega_{{\rm l},q}}}\frac{1}{(2\pi)^{2}}\left(a_{\mathbf{q}}+a_{-\mathbf{q}}^{\dagger}\right)\mathbf{e}_{\mathbf{q}},$ (S45)
@@
$\varphi^{\rm 2D}(\mathbf{r})\equiv\varphi(\mathbf{r},0)$ $=-\frac{i\gamma}{2\varepsilon_{0}}\int e^{i\mathbf{q}\cdot\mathbf{r}}\frac{1}{q}\frac{1}{\varepsilon+r_{\infty}q}\mathbf{q}\cdot\sqrt{\frac{\hbar A}{2\omega_{{\rm l},q}}}\frac{1}{(2\pi)^{2}}\left(a_{\mathbf{q}}+a_{-\mathbf{q}}^{\dagger}\right)\mathbf{e}_{\mathbf{q}}{\rm d}^{2}q$
$=-i\omega_{\rm t}\sum_{q}\sqrt{\frac{\hbar(r_{0}-r_{\infty})}{4\varepsilon_{0}A\omega_{{\rm l},q}}}\frac{1}{\varepsilon+r_{\infty}q}\left(a_{\mathbf{q}}e^{i\mathbf{q}\cdot\mathbf{r}}-a_{\mathbf{q}}^{\dagger}e^{-i\mathbf{q}\cdot\mathbf{r}}\right).$ (S46)